In the equation by - 3x = 6, for what value of b does the graph of the equation have a slope of 3?

Answer by ewatrrr(24785)

 
Hi,
In the equation   by+-+3x+=+6, 
for what value of b does the graph of the equation have a slope of 3?

 y = (3x+ 6)/ b  0r  y = (3/b)x + 6/b

 Slope = 3  ⇒  (3/b) = 3  0r b = 3/3  = 1


   When b = 1  the graph of the equation has a slope of 3

   y = 3x + 6



Wish You the Best in your Studies.